I'm trying to simplify this complex rational expression and I keep getting -3/x-y the answer key keeps saying it should be 3/x+y can you help?
3/y - 3/x ÷ x^2-y^2/xy=

Answer by John10(297)

(3/y - 3/x)/ [(x^2 - y^2)/ xy]
Work on the numerator first
3/y - 3/x = (3x - 3y)/xy = 3(x - y)/(xy)
The denominator:
[(x - y)(x + y)/(xy)]
Put them together in fraction form
[3(x - y)/(xy)] * [(xy)/(x - y)(x + y)]
Cancel like terms from the numerator and denominator
The result is:
3/(x + y) <=== answer key
Hope it helps you!
Online math tutor (john100185@yahoo.com)